### v2

@@ 1.1 Notation
    For given a matrix $M,$ $C(M)$ and $R(M)$ denote the column and row space of $M$, respectively. Also, $N(M)$ denotes the (right) null space of $M$.
@@ 2.1 Kronecker-Core Decomposition
In this section, we review the Kronecker-core decomposition proposed by [hoff2023]. Suppose $\Sigma\in\mathcal{S}_{p}^{++}$ and define a function $d:\mathcal{S}_{p_{1},p_{2}}^{++}\rightarrow\real$ by
(2) $d(K|\Sigma):=d(K_{2}\otimes K_{1}|\Sigma)=\textsf{tr}\left(\Sigma K^{-1}\right)+p_{1}\log|K_{2}|+p_{2}\log|K_{1}|,$
@@
Thus, $k(\Sigma)$ is the Kronecker maximum likelihood estimate (MLE) of $\Sigma$, representing the most separable component of $\Sigma$. Note that $k(\Sigma)$ uniquely exists for any $\Sigma\in\mathcal{S}_{p}^{++}$ [srivastava2008, hoff2023] and we refer the map $k:\mathcal{S}_{p}^{++}$ to as the Kronecker map.
To define the core component, let $h$ be a bijective square root map defined on $\mathcal{S}_{p_{1},p_{2}}^{++}$, e.g., a symmetric square root (i.e., $h(\mathcal{S}_{p_{1},p_{2}}^{++})\equiv\mathcal{S}_{p_{1},p_{2}}^{++}$) or a Cholesky factor (i.e., $h(\mathcal{S}_{p_{1},p_{2}}^{++})\equiv\mathcal{L}_{p_{1},p_{2}}^{++}$). By a slight abuse of notation, we write $h\in\mathcal{S}_{p_{1},p_{2}}^{++}$ (resp. $\mathcal{L}_{p_{1},p_{2}}^{++})$ when $h$ is taken to be the symmetric square root (resp. Cholesky factor). With a fixed choice of $h$, the core component of $\Sigma$ is defined to be $c(\Sigma)\equiv h(k(\Sigma))^{-1}\Sigma h(k(\Sigma))^{-\top}$. By the definition of the Kronecker map, it holds that $k(G\Sigma G^{\top})=Gk(\Sigma)G^{\top}$ for any $G\in GL_{p_{1},p_{2}}$ ([hoff2023], Proposition $2$). Thus, $k(c(\Sigma))=I_{p}$; that is, the Kronecker MLE of any core is $I_{p}$. This leads to the definition of the core covariance matrix as a positive definite matrix whose Kronecker MLE equals $I_{p}$. That is, the set of such covariance matrices is equivalent to $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv\left\{C\in\mathcal{S}_{p}^{++}:k(C)=I_{p}\right\}$. By the uniqueness of the Kronecker MLE and the bijectivity of the map $h$, the core component is uniquely defined for any $\Sigma$. Consequently, the map $c:\Sigma\in\mathcal{S}_{p}^{++}\rightarrow h(k(\Sigma))^{-1}\Sigma h(k(\Sigma))^{-\top}\in\mathcal{C}_{p_{1},p_{2}}^{++}$ is well-defined and referred to as the core map. By the definition of the maps $k$ and $c$, every $\Sigma$ admits a unique and identifiable Kronecker-core decomposition (KCD) as $h(k(\Sigma))c(\Sigma)h(k(\Sigma))^{\top}$ (see Proposition $5$ of [hoff2023]). The definitions of the separable and core components are summarized below.
@@
By the construction, $c(\Sigma)=I_{p}$ if and only if $\Sigma=k(\Sigma)\in\mathcal{S}_{p_{1},p_{2}}^{++}$. Namely, the only separable core is $I_{p}$. As discussed in Section 1, the Kronecker MLE $K$ may also uniquely exist for $\Omega\in\mathcal{S}_{p,r}^{+}$ if $r>p_{1}/p_{2}+p_{2}/p_{1}$ [drton2021, derksen2021, soloveychik2016] by taking $\Sigma=\Omega$ in (2). Provided that $K$ uniquely exists, its core component $C$ can be also uniquely defined by whitening $\Omega$ via $K^{1/2}$ as above. Since $K^{1/2}$ is non-singular, $\Omega$ and $C$ shares the same rank as $r$. Suppose $C=AA^{\top}$ for $A=[\text{vec}(A_{1}),\ldots,\text{vec}(A_{r})]\in\real_{*}^{p\times r}$ with $A_{i}\in\real^{p_{1}\times p_{2}}$. By Proposition $3$ of [hoff2023], $\tilde{A}:=(A_{1},\ldots,A_{r})$ should satisfy that
(3) $\textsf{tr}_{1}(C)\equiv\sum_{i=1}^{r}A_{i}A_{i}^{\top}=p_{2}I_{p_{1}},\quad\textsf{tr}_{2}(C)\equiv\sum_{i=1}^{r}A_{i}^{\top}A_{i}=p_{1}I_{p_{2}}.$
This motivates the set of the rankr-r core covariance matrices defined as $\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}\equiv\left\{C\in\mathcal{S}_{p,r}^{+}:\textsf{tr}_{1}(C)=p_{2}I_{p_{1}},\textsf{tr}_{2}(C)=p_{1}I_{p_{2}}\right\}$. Note that if $r=p$, $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}$.
We shall connect rankr-r cores to statistical applications, thereby motivating the study of rankr-r cores. Observe that the set $\mathcal{C}_{p_{1},p_{2}}^{++}$ is convex by (3). Using this convexity, [hoff2023] proposed a core shrinkage estimator (CSE) that shrinks the sample core toward the unique separable core, $I_{p}$. However, if the population core exhibits a low-dimensional feature, the CSE can be subject to over-parameterization when $n<p$. Specifically, [hoff2023] discussed a partial-isotropy structure as a possible structural assumption on $c(\Sigma)$, following the approach in factor analysis. Namely, $c(\Sigma)$ is represented as $BB^{\top}+\lambda I_{p}$ for some $B\in\real_{*}^{p\times r}$ and $\lambda>0$. By the construction of CSE, its partial-isotropy rank is $n$ when $n<p$ (see Section 3.13.13.23.2 of [hoff2023]), which is typically larger than $r$. Thus, it may over-parameterize such a $c(\Sigma)$. Nevertheless, they did not pursue incorporating the partial-isotropy structure of $c(\Sigma)$ in estimation themselves due to a lack of understanding of such a core. By the linear system in (2), there are constraints on $B$ and $\lambda$, compared to a usual partial-isotropy covariance. The following implies that a partial-isotropy rankr-r core is a non-trivial convex combination of a rankr-r core and a trivial core $I_{p}$, leading to the study of rankr-r cores.
Proposition 1.
@@
Since $BB^{\top}$ is positive semi-definite, so are its partial traces [zhang2012, zhang2011]. Hence, we should have that $\lambda\leq 1$. Note that the linear system in (3) implies that $\textsf{tr}(C)=p$. Therefore, if $\lambda=1$, $\textsf{tr}(BB^{\top})=0$ so that $BB^{\top}=\mathbf{0}_{p\times p}$, contradicting the assumption that $B\in\real_{*}^{p\times r}$. Thus, $\lambda<1$. Parameterizing $BB^{\top}$ by $(1-\lambda)AA^{\top}$ for some $A\in\real_{*}^{p\times r}$, we have that
@@
implying that $AA^{\top}$ is a rankr-r core.
@@
The reason is that the threshold on $r$ for which the Kronecker MLE exists is understood in a generic (almost sure) sense [derksen2021]. In a strict algebra sense, however, the Kronecker MLE may not exist for rankr-r $\Omega$ even if $r$ satisfies the threshold as seen above. Furthermore, unless $r=p$, $\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}$ may not be a smooth manifold. In Section 4.1, we show that the singularity preventing $\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}$ from being a manifold in view of Sard’s theorem ([lee2012], Theorem $6.10$) corresponds to the set of canonically decomposable matrices. Here, $(A_{1},\ldots,A_{r})\in(\real^{p_{1}\times p_{2}})^{r}$ is canonically decomposable if there exists $(P,Q)\in GL_{p_{1}}\times GL_{p_{2}}$ such that $PA_{i}Q^{-1}$ is of a non-trivial block-diagonal form for each $i\in[r]$. After removing this set from $\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}$, the remaining set $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a smooth manifold, as shown in Section 4.2.
@@ 2.2 Riemannian Manifolds
In this section, we briefly review some geometric properties of Riemannian manifolds. For details, we refer the reader to [lee2012, lee2018, absil2008, boumal2023]. Suppose $(\mathcal{M},g)$ is a Riemannian manifold, where $\mathcal{M}$ is a smooth manifold equipped with a Riemannian metric $g$. The Riemannian metric $g:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\real$ defines an inner product on each tangent space $T_{x}\mathcal{M}$, varying smoothly with $x\in\mathcal{M}$. A smooth curve $\gamma_{x}^{v}:[0,1]\rightarrow\mathcal{M}$ emanating from $x\in\mathcal{M}$ in the direction of $v\in T_{x}\mathcal{M}$ is geodesic, i.e., a locally shortest curve with zero acceleration. Then the exponential map $\operatorname{Exp}_{x}:T_{x}\mathcal{M}\rightarrow\mathcal{M}$ is defined by $\operatorname{Exp}_{x}(v):=\gamma_{x}^{v}(1)$. Suppose $f$ is a smooth function on $\mathcal{M}$. The Riemannian gradient $\operatorname{grad}f(x)$ of $f$ at $x\in\mathcal{M}$ is the unique tangent vector in $T_{x}\mathcal{M}$ satisfying that for any $v\in T_{x}\mathcal{M}$,
@@
where $D_{v}f(x)$ is a directional derivative of $f(x)$ along $v$. The Riemannian Hessian operator of $f$, denoted $\operatorname{Hess}f(x):T_{x}\mathcal{M}\rightarrow T_{x}\mathcal{M}$, is then defined to be a covariant derivative of the Riemannian gradient. By (5.35)–(5.36) of [absil2008], for a geodesic $\gamma_{x}^{v}$,
@@
Since $\operatorname{Hess}f(x)[\cdot,\cdot]$ is a symmetric bilinear form on $T_{x}\mathcal{M}$ (see (5.31) of [absil2008]), the polarization identity implies that
(4) $g_{x}(\operatorname{Hess}f(x)[v],w)=\frac{\operatorname{Hess}f(x)[v+w,v+w]-\operatorname{Hess}f(x)[v,v]-\operatorname{Hess}f(x)[w,w]}{2},$
@@ 2.2.1 Riemannian Geometry of $\mathcal{S}_{p}^{++}$ and $\mathbb{P}(\mathcal{S}_{p}^{++})$
We review the Riemannian geometry of $\mathcal{S}_{p}^{++}$ and $\mathbb{P}(\mathcal{S}_{p}^{++})$ under the affine-invariant metric $g^{\text{AI}}$ [skovgaard1984, pennec2006, moakher2005]. The tangent spaces of each manifold are given by
@@
The smooth manifold $\mathbb{P}(\mathcal{S}_{p}^{++})$ is a totally geodesic submanifold of $\mathcal{S}_{p}^{++}$ under $g^{\text{AI}}$. Also, the orthogonal projection of $V\in T_{\Sigma}\mathcal{S}_{p}^{++}$ onto $T_{\Sigma}\mathbb{P}(\mathcal{S}_{p}^{++})$ (see ($32$) of [simonis2025]) is given by
(5) $\mathcal{P}_{\Sigma}(V):=V-\textsf{tr}\left(\Sigma^{-1}V\right)\Sigma/p.$

### v3

@@ 1.1 Notations
    For given a matrix $M,$ $C(M)$ and $N(M)$ denote the column and (right) null space of $M$, respectively.
@@ 2.1 Kronecker-core decomposition
In this section, we review the Kronecker-core decomposition proposed by [33]. Suppose $\Sigma\in\mathcal{S}_{p}^{++}$ and define a function $d:\mathcal{S}_{p_{1},p_{2}}^{++}\rightarrow\real$ by
$d(K|\Sigma):=d(K_{2}\otimes K_{1}|\Sigma)=\textsf{tr}\left(\Sigma K^{-1}\right)+p_{1}\log|K_{2}|+p_{2}\log|K_{1}|,$ (3)
@@
Thus, $k(\Sigma)$ is the Kronecker maximum likelihood estimate (MLE) of $\Sigma$, representing the most separable component of $\Sigma$. Note that $k(\Sigma)$ uniquely exists for any $\Sigma\in\mathcal{S}_{p}^{++}$ [62, 33] and we refer the map $k:\mathcal{S}_{p}^{++}$ to as the Kronecker map.
To define the core component, let $h$ be a bijective square root map defined on $\mathcal{S}_{p_{1},p_{2}}^{++}$, e.g., a symmetric square root (i.e., $h(\mathcal{S}_{p_{1},p_{2}}^{++})\equiv\mathcal{S}_{p_{1},p_{2}}^{++}$) or a Cholesky factor (i.e., $h(\mathcal{S}_{p_{1},p_{2}}^{++})\equiv\mathcal{L}_{p_{1},p_{2}}^{++}$). By a slight abuse of notation, we write $h\in\mathcal{S}_{p_{1},p_{2}}^{++}$ (resp. $\mathcal{L}_{p_{1},p_{2}}^{++})$ when $h$ is taken to be the symmetric square root (resp. Cholesky factor). With a fixed choice of $h$, the core component of $\Sigma$ is defined to be $c(\Sigma)\equiv h(k(\Sigma))^{-1}\Sigma h(k(\Sigma))^{-\top}$. By the definition of the Kronecker map, it holds that $k(G\Sigma G^{\top})=Gk(\Sigma)G^{\top}$ for any $G\in GL_{p_{1},p_{2}}$ ([33], Proposition $2$). Thus, $k(c(\Sigma))=I_{p}$; that is, the Kronecker MLE of any core is $I_{p}$. This leads to the definition of the core covariance matrix as a positive definite matrix whose Kronecker MLE equals $I_{p}$. That is, the set of such covariance matrices is equivalent to $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv\left\{C\in\mathcal{S}_{p}^{++}:k(C)=I_{p}\right\}$. By the uniqueness of the Kronecker MLE and the bijectivity of the map $h$, the core component is uniquely defined for any $\Sigma$. Consequently, the map $c:\Sigma\in\mathcal{S}_{p}^{++}\rightarrow h(k(\Sigma))^{-1}\Sigma h(k(\Sigma))^{-\top}\in\mathcal{C}_{p_{1},p_{2}}^{++}$ is well-defined and referred to as the core map. By the definition of the maps $k$ and $c$, every $\Sigma$ admits a unique and identifiable Kronecker-core decomposition (KCD) as $h(k(\Sigma))c(\Sigma)h(k(\Sigma))^{\top}$ (see Proposition $5$ of [33]). The definitions of the separable and core components are summarized below.
@@
By the construction, $c(\Sigma)=I_{p}$ if and only if $\Sigma=k(\Sigma)\in\mathcal{S}_{p_{1},p_{2}}^{++}$. Namely, the only separable core is $I_{p}$. As discussed in Section 1, the Kronecker MLE $K$ may also uniquely exist for $\Omega\in\mathcal{S}_{p,r}^{+}$ if $r>p_{1}/p_{2}+p_{2}/p_{1}$ [23, 19, 61] by taking $\Sigma=\Omega$ in (3). Provided that $K$ uniquely exists, its core component $C$ can be also uniquely defined by whitening $\Omega$ via $K^{1/2}$ as above. Since $K^{1/2}$ is non-singular, $\Omega$ and $C$ shares the same rank as $r$. Suppose $C=AA^{\top}$ for $A=[\text{vec}(A_{1}),\ldots,\text{vec}(A_{r})]\in\real_{*}^{p\times r}$ with $A_{i}\in\real^{p_{1}\times p_{2}}$. By Proposition $3$ of [33], $\tilde{A}:=(A_{1},\ldots,A_{r})$ should satisfy that
$\textsf{tr}_{1}(C)\equiv\sum_{i=1}^{r}A_{i}A_{i}^{\top}=p_{2}I_{p_{1}},\quad\textsf{tr}_{2}(C)\equiv\sum_{i=1}^{r}A_{i}^{\top}A_{i}=p_{1}I_{p_{2}}.$ (4)
This motivates the set of the rankr-r core covariance matrices defined as
$\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}\equiv\left\{C\in\mathcal{S}_{p,r}^{+}:\textsf{tr}_{1}(C)=p_{2}I_{p_{1}},\textsf{tr}_{2}(C)=p_{1}I_{p_{2}}\right\}.$
Note that if $r=p$, $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}$.
We shall connect rankr-r cores to statistical applications, thereby motivating the study of rankr-r cores. Observe that the set $\mathcal{C}_{p_{1},p_{2}}^{++}$ is convex by (4). Using this convexity, [33] proposed a core shrinkage estimator (CSE) that shrinks the sample core toward the unique separable core, $I_{p}$. However, if the population core exhibits a low-dimensional feature, the CSE can be subject to over-parameterization when $n<p$. Specifically, [33] discussed a partial-isotropy structure as a possible structural assumption on $c(\Sigma)$, following the approach in factor analysis. Namely, $c(\Sigma)$ is represented as $BB^{\top}+\lambda I_{p}$ for some $B\in\real_{*}^{p\times r}$ and $\lambda>0$. By the construction of CSE, its partial-isotropy rank is $n$ when $n<p$ (see Section 3.13.13.23.2 of [33]), which is typically larger than $r$. Thus, it may over-parameterize such a $c(\Sigma)$. Nevertheless, they did not pursue incorporating the partial-isotropy structure of $c(\Sigma)$ in estimation themselves due to a lack of understanding of such a core. By the linear system in (3), there are constraints on $B$ and $\lambda$, compared to a usual partial-isotropy covariance. The following implies that a partial-isotropy rankr-r core is a non-trivial convex combination of a rankr-r core and a trivial core $I_{p}$, leading to the study of rankr-r cores.
Proposition 2.2.
@@
Since $BB^{\top}$ is positive semi-definite, so are its partial traces [74, 73]. Hence, we should have that $\lambda\leq 1$. Note that the linear system in (4) implies that $\textsf{tr}(C)=p$. Therefore, if $\lambda=1$, $\textsf{tr}(BB^{\top})=0$ so that $BB^{\top}=\mathbf{0}_{p\times p}$, contradicting the assumption that $B\in\real_{*}^{p\times r}$. Thus, $\lambda<1$. Parameterizing $BB^{\top}$ by $(1-\lambda)AA^{\top}$ for some $A\in\real_{*}^{p\times r}$, we have that
@@
implying that $AA^{\top}$ is a rankr-r core. ∎
Remark 1.
Regarding the condition on $r$ in Proposition 2.2, recall that it arises from the sample size threshold for which the Kronecker MLE exists. In fact, by Theorem $1.2$ of [19], the other scenarios on $(p_{1},p_{2},r)$ that admit the Kronecker MLE are either $p_{1}^{2}+p_{2}^{2}-rp_{1}p_{2}=0$, which is equivalent to $(p_{1},p_{2},r)=(p_{1},p_{1},2)$, or $p_{1}^{2}+p_{2}^{2}-rp_{1}p_{2}=d^{2}$ for $d=\text{gcd}(p_{1},p_{2})$. Assuming $p_{1}\geq p_{2}$ without loss of generality, it can be shown that the latter holds only when $(p_{1},p_{2},r)=(p_{2}r,p_{2},r),((k+1)m,km,2)$, $(p_{2}^{2}-1,p_{2},p_{2})$ for any $k,m,p_{2},r\in\mathbb{N}$ (see [63] also). Compared to the regime where $p_{1}/p_{2}+p_{2}/p_{1}<r$, equivalently $p_{1}^{2}+p_{2}^{2}-rp_{1}p_{2}<0$, the other scenarios are highly restrictive, as generic $(p_{1},p_{2},r)$ do not satisfy them. On the other hand, the regime where $p_{1}/p_{2}+p_{2}/p_{1}<r$ applies to generic $(p_{1},p_{2},r)$ and allows freeness in the choice of $r$.
@@
The reason is that the threshold on $r$ for which the Kronecker MLE exists is understood in a generic (almost sure) sense [19]. In a strict algebra sense, however, the Kronecker MLE may not exist for rankr-r $\Omega$ even if $r$ satisfies the threshold as seen above. Furthermore, unless $r=p$, $\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}$ may not be a smooth manifold. In Section 4.1, we show that the singularity preventing $\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}$ from being a manifold in view of Sard’s theorem ([38], Theorem $6.10$) corresponds to the set of canonically decomposable matrices. Here, $(A_{1},\ldots,A_{r})\in(\real^{p_{1}\times p_{2}})^{r}$ is canonically decomposable if there exists $(P,Q)\in GL_{p_{1}}\times GL_{p_{2}}$ such that $PA_{i}Q^{-1}$ is of a non-trivial block-diagonal form for each $i\in[r]$. After removing this set from $\tilde{\mathcal{C}}_{p_{1},p_{2},r}^{+}$, the remaining set $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a smooth manifold, as shown in Section 4.2.
@@ 2.2 Riemannian manifolds
In this section, we briefly review some geometric properties of Riemannian manifolds. For details, we refer the reader to [38, 39, 1, 9]. Suppose $(\mathcal{M},g)$ is a Riemannian manifold, where $\mathcal{M}$ is a smooth manifold equipped with a Riemannian metric $g$. The Riemannian metric $g:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\real$ defines an inner product on each tangent space $T_{x}\mathcal{M}$, varying smoothly with $x\in\mathcal{M}$. A smooth curve $\gamma_{x}^{v}:[0,1]\rightarrow\mathcal{M}$ emanating from $x\in\mathcal{M}$ in the direction of $v\in T_{x}\mathcal{M}$ is geodesic, i.e., a locally shortest curve with zero acceleration. Then the exponential map $\operatorname{Exp}_{x}:T_{x}\mathcal{M}\rightarrow\mathcal{M}$ is defined by $\operatorname{Exp}_{x}(v):=\gamma_{x}^{v}(1)$. Suppose $f$ is a smooth function on $\mathcal{M}$. The Riemannian gradient $\operatorname{grad}f(x)$ of $f$ at $x\in\mathcal{M}$ is the unique tangent vector in $T_{x}\mathcal{M}$ satisfying that for any $v\in T_{x}\mathcal{M}$,
@@
where $D_{v}f(x)$ is a directional derivative of $f(x)$ along $v$. The Riemannian Hessian operator of $f$, denoted $\operatorname{Hess}f(x):T_{x}\mathcal{M}\rightarrow T_{x}\mathcal{M}$, is then defined to be a covariant derivative of the Riemannian gradient. By (5.35)–(5.36) of [1], for a geodesic $\gamma_{x}^{v}$,
@@
Since $\operatorname{Hess}f(x)[\cdot,\cdot]$ is a symmetric bilinear form on $T_{x}\mathcal{M}$ (see (5.31) of [1]), the polarization identity implies that
$g_{x}(\operatorname{Hess}f(x)[v],w)=\frac{\operatorname{Hess}f(x)[v+w,v+w]-\operatorname{Hess}f(x)[v,v]-\operatorname{Hess}f(x)[w,w]}{2},$ (5)
@@ 2.2.1 Riemannian geometry of $\mathcal{S}_{p}^{++}$ and $\mathbb{P}(\mathcal{S}_{p}^{++})$
We review the Riemannian geometry of $\mathcal{S}_{p}^{++}$ and $\mathbb{P}(\mathcal{S}_{p}^{++})$ under the affine-invariant metric $g^{\text{AI}}$ [60, 54, 48]. The tangent spaces of each manifold are given by
@@
The smooth manifold $\mathbb{P}(\mathcal{S}_{p}^{++})$ is a totally geodesic submanifold of $\mathcal{S}_{p}^{++}$ under $g^{\text{AI}}$. Also, the orthogonal projection of $V\in T_{\Sigma}\mathcal{S}_{p}^{++}$ onto $T_{\Sigma}\mathbb{P}(\mathcal{S}_{p}^{++})$ (see ($32$) of [59]) is given by
$\mathcal{P}_{\Sigma}(V):=V-\textsf{tr}\left(\Sigma^{-1}V\right)\Sigma/p.$ (6)